Mia’s Animal Shelter performs three services: housing and finding homes for stray and unwanted animals, providing health care and neutering services for the animals, and pet training services. One facility is dedicated to housing animals waiting to be adopted. A second facility houses veterinarian services. A third facility houses the director, administration staff, and several dog trainers. This facility also has several large meeting rooms that are frequently used for classes conducted by animal trainers. The trainers work with all of the shelter’s animals to ensure that they are relatively easy to manage. They also provide dog obedience classes for families who adopt animals.

Estimated annual costs for the animal shelter and its services are as follows: ( in $ )

Director and administration staff salaries : 60,000

Housing service’s employees’ salaries : 100,000

Veterinarians and technicians’ salaries : 150,000

Animal trainers’ salaries : 40,000

Food and supplies : 125,000

Building-related costs : 200,000

On average, 75 animals per day are housed at the facility, or about 27,375 (75 × 365) animal days in total. In addition, the trainers offer about 125 classes for about 30 weeks throughout the year. On average, 10 families attend each class. Last year the veterinarian clinic experienced 5,000 animal visits.

One of the administration staff, Eden, just graduated from an accounting program and would like to set up an ABC system for the shelter so that the director can better understand thecost for each of the shelter’s services. He gathers the following information:

Square metres for each facility:

Housing service : 5,000 sq meters
Director and training : 3,000 sq meters

Veterinarian clinic : 2,000 square meters

Percentage of trainer time used in classes : 50%

Supplies used for veterinarian services : $75,000

Question 1: ABC

  1. Identify three cost pools and assign costs to them, considering the three cost objects of interest. Determine an activity driver for each cost pool identified. Explain your choice. (10 marks)

  2. Using the activity drivers you identified in part (i) calculate the direct cost per unit of activity driver for each pool and explain what the cost per unit of activity driver for each pool means. Show ALL your workings. (10 marks)
Answer #1

ANSWER

1) Cost driver for each cost pool :

Direct and staff salaries                             $60,000 - Animal days

Animal shelter employees’ salaries          100,000 - Number of animal housed

Veterinarians and technicians                   150,000 - animals visited

Animal trainers                                         40,000 - Per training class

Food and supplies                                     125,000 - Number of animal housed

Building-related costs                               200,000 - Square footage occupied

2) Allocation rates :

Direct and staff salaries $60,000 - Animal days = $60000/27375 = $2.20 per animal day

Animal shelter employees’ salaries 100,000 - Number of animal housed = 100000/4500 = $22.22 / animal housed.

Veterinarians and technicians 150,000 - animals visited = 150000 / 5000 = $30 per animal visit

Animal trainers 40,000 - Per training class = $40000 / 125 = $320 / training class

Food and supplies 125,000 - Number of animal housed = 125000 / 4500 = $27.78 / animal housed

Building-related costs 200,000 - Square footage occupied = 200000 / 10000 = $20 / Sq. footage

3) The advantages of ABC system are : specific activity cost rate, proper allocation of cost, accurate costing of job, contract costing quatation become easy.

The disadvantages of ABC system are : time consuming and costly method, huge difference between cost drivers, disproportionate allocation of costs.

------------------------------------------------------

*****************THANK YOU******************